where do the two lines cross?
y=2x-6
3y-5x=2

Answer by checkley77(12569)

y=2x-6
y-2x=-6 multiply by -3 & add.
3y-5x=2
-3y+6x+18
-------------
x=20 ans.
y=2*20=-6
y=40-6
y=34 ans.
Proof:
3*34-5*20=2
102-100=2
2=2